OpenStudy (anonymous):

Can anyone help with evaluating a series? I first answered it via proof by induction but I'm sure there must be a way to evaluate it.

OpenStudy (anonymous):

Here's the question: Show that the sum S(N) of the first N terms of the series \[\frac{ 1 }{ 1\times2\times3 } + \frac{ 3 }{ 2\times3\times4 } + \frac{ 5 }{ 3\times4\times5 } + ... + \frac{ 2n-1 }{n(n+1)(n+2)} + ...\] is \[\frac{ 1 }{ 2 }(\frac{ 3 }{ 2 } + \frac{ 1 }{ N+1 } - \frac{5}{N+2})\]

OpenStudy (anonymous):

I'm not exactly certain of the Solution. Though I did manage to find a simplified form of Nth term. T(N) = \[\frac{ 2n-1 }{ (n)(n+1)(n+2) } = \frac{ A }{ n } + \frac{ B }{( n+1) } + \frac{ C }{ n+2 }\] I'm trusting you can do the simplification This implies 2n-1 = A(n+1)(n+2)+B(n)(n+2) +C (n)(n+1) =(A+B+C)n^{2} + (3A+ 2B+ C)n + 2A Which gives us, A+B+C=0 3A+2B+C=2 which implies 2A +B=2 2A=-1 This means A=-1/2, B=3, C= 5/2 Thus the n'th Term becomes \[T(n) = \frac{ -1 }{ 2n } + \frac{ 3 }{ n+1 } + \frac{ 5 }{ 2(n+2) }\] Can't really get beyond this, though.

OpenStudy (anonymous):

And here's a correction, it's -5, not +5... -_-"

OpenStudy (cwrw238):

If I remember correctly this type of series can be solved by the method of differences I'd have to look it up before I am sure. if I have time I'll get back to you today.

OpenStudy (anonymous):

Thanks for the help - I did try to simplify the nth term and got what you did, but I have no idea how to sum the terms. I'll see if I can find anything about the method of differences with a quick google. :P